determine the general/particular solution for each equation using the applicable solution to equations of order one (separable, homogenous,linear,exact)


3x(xy-2)dx + (x^3=2y)dy = 0


1
Expert's answer
2022-02-14T10:48:00-0500

Let us determine the general solution of the differential equation


"3x(xy-2)dx + (x^3-2y)dy = 0."


Taking into account that


"\frac{\partial( 3x(xy-2))}{\partial y} =3x^2=\frac{\partial (x^3-2y)}{\partial x},"


we conclude that this equation is exact, and hence there is a function "u=u(x,y)" such that


"\frac{\partial u}{\partial x}=3x(xy-2)=3x^2y-6x,\ \ 
\frac{\partial u}{\partial y}=x^3-2y."


It follows that "u=x^3y-3x^2+c(y)," and thus


"\frac{\partial u}{\partial y}=x^3+c'(y)=x^3-2y."


Therefore, "c'(y)=-2y," and thus "c(y)=-y^2+C."


We conclude that the general solution is of the differential equation


"x^3y-3x^2-y^2=C."